OpenStudy (anonymous):

The formula s = p/6 gives the side length s of a regular hexagon with perimeter P. What is the side length of a regular hexagon with perimeter 15 inches? A. 2.5 in. B. 9 in. C. 21 in. D. 90 in.

OpenStudy (drummergirl3):

15/6.... sides of hexagon are 6... so solve..

OpenStudy (drummergirl3):

A

OpenStudy (jhulian):

s=p/6 p=15 s=15/6 s=2.5

OpenStudy (anonymous):

2.5?

OpenStudy (anonymous):

Thank you!

OpenStudy (drummergirl3):

No Problem! cute owl btw ;D